OpenStudy (anonymous):

Cos theta= 2/7, Cot theta less than 0, find Sin theta

terenzreignz (terenzreignz):

In which quadrant would this lie?

OpenStudy (anonymous):

None. It's suppose to be less than 0. Meaning negative.

terenzreignz (terenzreignz):

What is? Sine? Yes, Sine would be negative. There are two main ways to go about this, by drawing a triangle or just using identities. Which one are you more comfortable with?

OpenStudy (anonymous):

I think I got it. I hope I am correct. Sin^2 + Cos^2= 1 Sin^2 + (2/7)^2 = Sin^2 + 4/49= Sin^2= 1/49-4/49 Sin^2=45/49 √Sin^2= √45/7 Sin=3√5/7 Is that correct?